OpenStudy (anonymous):

If 2.6 plus 9.3 times a number equals the sum of the number and 40.3 what is the number?

OpenStudy (cwrw238):

2.6 + 9.3x = x + 40.3 solve for x

OpenStudy (cwrw238):

8.3x = 40.3 - 2.6 = 37.7 x = 37.7 / 8.3 = 4.54 to 2 dec places

OpenStudy (anonymous):

oh ok i think i got it